3. Gizmo, now a not-so-good Nascar driver, keeps track of his speedometer every minute. The table below shows what he found, PLEASE SHOW STEPS IF POSSIBLE
Answer #1

We know that total distance = Avg speed*time

Avg speed = (55+57+60+58+60+62+57+61+64+59)/10 = 59.3mph

Total time = 10minutes = 10/60 hour

Total distance = 59.3*10/60 = 9.88 miles